OpenStudy (anonymous):

Write an equation of a line perpendicular to y=1/2x-5 in slope-intercept form that passes through the points (2,-7). A.) y=1/2x-7 B.) y=2x+2 C.) y=-2x-5 D.) y=-2x-3

OpenStudy (anonymous):

Please explain to!(:

OpenStudy (***[isuru]***):

ok... there r some basic things u need to know first... if there r 2 lines which r perpendicular to each other... lets say the slope of one line is "m" then the slope of the other line is "(-1/m) for an example if the slope of line is "4" then the slope of the line perpendicular to it is "(-1/4)" so... first of all what is the slope of y=1/2x-5 ? can u find it ?

OpenStudy (anonymous):

1/2?

OpenStudy (***[isuru]***):

yep... so what will be the slope of a line perpendicular to it ?

OpenStudy (anonymous):

-1/1/2?

OpenStudy (***[isuru]***):

yep... now if u simplify it u will get -2.. got it ?

OpenStudy (anonymous):

yes

OpenStudy (***[isuru]***):

and now... the general format of a straight line is y = mx + b where m = slope b = y - intercept.. we already know the slope of the line we want .. so we can write it as y= -2x + b now all we have to do is find b to do that we can use the given pair of coordinates (2 , -7) since the line pass through this point... we can plug these values in to our equation.. in ( 2 , -7) x value = 2 y value = -7 so... if we plug them y = -2x + b -7 = -2(2) + b what will u get for b ?

OpenStudy (***[isuru]***):

- 7 = -4 + b b = -7 + 4 = -3 so... the equation of the line will be y = -2x - 3 hope this will help ya :)

OpenStudy (anonymous):

it helped a lot! thank you so much!